OpenStudy (anonymous):

The total rainfall in Forks, Washington, this October was in. In September, 60% less rain fell than in October. How much rain fell in September in Forks, Washington? Express your answer as a mixed number in simplest form.

OpenStudy (anonymous):

@acxbox22

OpenStudy (anonymous):

@HelpBlahBlahBlah

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@brittanydosey

OpenStudy (anonymous):

i am so confused... ;-;

OpenStudy (brittanydosey):

what is the number?

OpenStudy (anonymous):

oh sorry, 8 3/8

OpenStudy (anonymous):

mixxed number

OpenStudy (acxbox22):

multiply 8 3/8 by 6/10 67/8 x 6/10= ?

OpenStudy (brittanydosey):

i believe that's correct but im still thinking

OpenStudy (anonymous):

i came out with, 402 0ver 80

OpenStudy (anonymous):

should i now divid my numbers?

OpenStudy (anonymous):

Nvm ill just google it or somthing.... sorry for the wast of time

OpenStudy (acxbox22):

oh wait it is 67/8 x 4/10 since it is 60 percent less 67/8 x 4/10 = 67/20 now convert that to a mixed number

OpenStudy (brittanydosey):

yeah thats what i was thinking but i didnt think it was correct.

OpenStudy (anonymous):

6 7/20?

OpenStudy (brittanydosey):

i think you are correct.. but i am not 100% sure.

OpenStudy (anonymous):

me neither :P

OpenStudy (anonymous):

idk who to medal though :P

OpenStudy (anonymous):

it goes to you lol

OpenStudy (brittanydosey):

thank you. (:

OpenStudy (anonymous):

np lol :3